# Instant Runoff Voting and the Reinforcement Paradox

### v1

@@ Abstract.
We analyze the susceptibility of instant runoff voting (IRV) to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two sets of ballots such that a given losing candidate wins each of the sub-elections. Using these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We also analyze the frequency with which the paradox in a large dataset of real-world ranked-choice elections to provide empirical probabilities. Our general finding is that IRV is highly susceptible to this paradox in three-candidate elections.
@@ 1. Introduction
The third strand is the expanding empirical literature on IRV, enabled by the increasing availability of ranked-choice ballot data from jurisdictions such as Australia [42], the United States [33], and Scotland [23]. Some of this literature concerns the empirical probability that IRV produces a particular kind of paradox. Graham-Squire and Zayatz [14], Graham-Squire and McCune [16], and McCune and Graham-Squire [23] estimate the frequency with which real-world IRV elections demonstrate monotonicity and no-show paradoxes. Other empirical work includes case studies of particularly interesting elections which demonstrate a range of different paradoxes; see [19, 24], for example. Beyond work on paradoxes, empirical research on IRV has analyzed rates of ballot errors in IRV elections [6, 34], the extent of ballot exhaustion [5, 16], the rate at which IRV exhibits the so-called “spoiler effect” [26], the extent to which ballot truncation can affect the IRV winner [8, 18, 43], and the frequency with which IRV selects the Condorcet winner [16, 41, 42]. There is also a growing, related empirical literature concerning the method of the single transferable vote (STV), a multiwinner version of IRV which is designed to achieve proportional representation given a set of preference ballots. This literature also investigates paradoxical and Condorcet-related outcomes [23, 25], as well as STV’s propensity to achieve proportional representation [1, 3, 29].
@@
Previous work on the reinforcement paradox is fairly limited. The notion of a reinforcement paradox dates to the work of Young [44, 45], who defined the concept and proved essentially that the only reasonable voting methods not susceptible to the paradox are positional scoring rules. Felsenthal and Nurmi [12] show that various Condorcet inconsistent voting rules (including IRV) are susceptible to the paradox even in “tame” profiles in which a Condorcet winner exists and the voting rule chooses this candidate. Brandt et al. [2] provide some “minimal” examples of the reinforcement paradox under IRV where the number of candidates and voters is small. We are aware of only two previous papers which attempt to measure the frequency with which IRV produces reinforcement paradoxes, both of which focus on the three-candidate case. Plassmann and Tideman [36] measure this frequency under a spatial model, and Courtin et al. [7] use the impartial culture and impartial anonymous culture models of voter behavior. Our work differs from these studies in significant aspects which we explicate more fully below. In brief, these two studies tend to generate two elections and then say that the reinforcement paradox has occurred if these two elections have the same winner but when the ballots from the two elections are combined the winner is different. By contrast, we consider only one election and say the paradox has occurred in this election if there exists a way to partition the ballots into two sub-elections where the winners of the sub-elections are the same but differ from the winner of the original election.
@@
The paper is structured as follows. Section 2 gives preliminaries, including formal definitions of IRV and the reinforcement paradox. In Section 3 we provide necessary and sufficient conditions for a three-candidate election to demonstrate a reinforcement paradox. Proposition 1 gives these conditions assuming all voters provide a complete ranking of the candidates while Proposition 2 gives conditions when some voters cast partial ballots. Section 4 considers the minimum percentage a losing candidate can earn and yet the election demonstrates a reinforcement paradox. Section 5 estimates the probability that an election demonstrates a paradox under various theoretical models of voter behavior. Section 6 analyzes real-world IRV elections from the United State and Scotland. This section provides empirical probabilities using a large dataset of elections, and provides a probabilistic analysis using non-parametric bootstrapping. Section 7 gives a discussion of our results, and Section 8 concludes.
@@ 3. Necessary and Sufficient Conditions
The point of this article is to investigate the probability with which a three-candidate election demonstrates a reinforcement paradox. To examine such probabilities, we first must know when an election demonstrates a paradox. This section characterizes when an election demonstrates a reinforcement paradox, giving sets of inequalities in terms of the parameters provided in Table 3 for the case of complete ballots or the parameters provided in Table 4 for the partial ballot case.
First, we note that if one of the candidates earns an initial majority of first-place votes (we term such a candidate a “majority candidate”) then the election cannot demonstrate a reinforcement paradox. That is, we cannot observe a paradox in the complete ballot case if $X_{1}+X_{2}>Y_{1}+Y_{2}+Z_{1}+Z_{2}$ or $Y_{1}+Y_{2}>X_{1}+X_{2}+Z_{1}+Z_{2}$ or $Z_{1}+Z_{2}>X_{1}+X_{2}+Y_{1}+Y_{2}$; similarly for the partial ballot case. The reason is that if a majority candidate exists, that candidate will earn a majority of first-place votes in either $\mathcal{B}_{1}$ or $\mathcal{B}_{2}$ for any partition of the ballot set.
@@
The conditions in Proposition 2 reduce to those in Proposition 1 when $X_{3}=Y_{3}=Z_{3}$ for both runner-up and plurality loser reinforcement paradoxes. The third condition $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})$ in Case (2) of Proposition 2 reduces to $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})$ and does not have an analogue in Proposition 1 because it is implied by the other conditions in Case (2) by the following argument. Note that
@@
Before sketching the proofs of the propositions, we note that runner-up and plurality loser reinforcement paradoxes can occur when the IRV and plurality winners are distinct or when they coincide, assuming either complete or partial ballots. In Case (2) of the runner-up paradox (plurality loser paradox respectively), it is easily seen that the runner-up (plurality loser respectively) wins both sub-elections only when different candidates are eliminated in the two sub-elections.
In Case (1) of the runner-up paradox, the runner-up must be the plurality winner, and the proof of the propositions show that if the two conditions are met, the runner-up wins a majority of first-place votes in one sub-election and wins in two rounds with the IRV winner being the plurality loser in the other sub-election. Note in Case (1), if $X_{1}>X_{2}$ then $Y^{\prime}>-X_{1}+X_{2}+Z^{\prime}$ is automatically satisfied. Thus, if among the voters whose first choice is the IRV winner, more voters second-rank the runner-up than second-rank the plurality loser, the paradox occurs whenever the runner up is the plurality winner ($Y^{\prime}>X^{\prime}$ or $Y>X$ in the complete ballot case).
@@
Candidate $C$ cannot have a majority of first-place votes in both subelections (otherwise it would have a majority in the original election). We claim that $C$ cannot have a majority of first-place votes in either subelection. To see this, suppose $C$ has a majority in $\mathcal{B}_{1}$. If $A$ is the plurality loser in $\mathcal{B}_{2}$, then
@@
Thus, $C$ must win in the second round after the elimination of a plurality loser in both subelections. It is easily seen that the plurality losers must be different in $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. So assume $B$ is the plurality loser in $\mathcal{B}_{1}$ and $A$ is the plurality loser in $\mathcal{B}_{2}.$ This yields the following set of inequalities.
@@
The inequalities created in each step of the iterative process in the proofs of the propositions can be used in reverse to identify election partitions leading to the paradoxes. For instance, from the proof of Case (1) of Proposition 1, we have
@@
Care must be taken when applying these bounds: the inequalities hold only in the limit as no allowance is made for the discrete mature of ballot counts and the possibility of ties. However, if the $X_{1},X_{2},$ etc. values in the original election are sufficiently distinct, these bounds can be used to identify two subelections.
For instance, in the complete ballot case, if $X_{1}=100,X_{2}=105,Y=215,Z_{1}=110$ and $Z_{2}=90$, these conditions allow us to select
$x_{1}=99,x_{2}=1,y=106\mbox{ and }z=197$
(among other possibilities), which leads to
@@
In real-world elections, the IRV runner-up tends to be a much stronger candidate than the plurality loser, and thus we expect that if an election demonstrates a plurality loser reinforcement paradox then the election will also demonstrate a runner-up paradox. However, the inequalities in Propositions 1 and 2 make clear this is not necessarily the case. Consider the profile shown in Table 5, for example. Neither Condition (1) nor (2) of Proposition 1 is satisfied, but the plurality loser inequalities are satisfied. Such preference profiles can arise when voter preferences are “single-peaked.” In this case, candidate $C$ is like a “centrist” candidate in a spatial model, as if $A$, $B$, and $C$ have been arranged on a left-to-right line, with $C$ in the middle. In Section 5 we use one-dimensional spatial models to investigate the frequency of reinforcement paradoxes when candidates can be arranged on a line.
Of course, real-world elections typically do not return ballot data resembling the profile in Table 5.
@@
1st $A$ $B$ $C$ $C$
2nd $C$ $C$ $A$ $B$
3rd $B$ $A$ $B$ $A$
Table 5. (Top) A profile which demonstrates a plurality loser, but not a runner-up, reinforcement paradox.
@@ 4. Threshold Results
Consider the profile in the top of Table 6, where $p_{A}$ is the proportion of voters who cast the ballot $A\succ B\succ C$ and also the proportion of voters who cast the ballot $A\succ C\succ B$, and $\epsilon$ and $\delta$ are positive and very small. Furthermore, assume $0.52p_{A}+\epsilon<$ 0.5 and $14p_{A}+\epsilon+\delta=$ 1 (which implies $\delta>\epsilon)$. Note that $A$ is the IRV winner and $C$ is the plurality loser.
This profile satisfies condition (2) of proposition 1 and thus the election demonstrates a runner-up reinforcement paradox. For completeness, we include a partition of the electorate into two sub-electorates in the bottom of Table 6. Let $\delta_{B}$ and $\delta_{A}$ be positive small parameters such that $20<\delta_{B}<\delta_{A}<\delta/$ 2. Note that $B$ wins when using the ballots of both sub-electorates separately. Since $\delta$ is arbitrary, the proposition is proven.∎
Num. Voters $p_{A}$ $p_{A}$ $2p_{A}+\epsilon$ $\delta$
1st $A$ $A$ $B$ $C$
2nd $B$ $C$ $A$ $A$
3rd $C$ $B$ $C$ $B$
$\delta-\delta_{A}$ $\delta-\delta_{B}$ $\delta$
$A$ $B$ $C$
$B$ $A$ $A$
$C$ $C$ $B$
$p_{A}-\delta+\delta_{A}$ $p_{A}$ $2p_{A}+\epsilon-\delta+\delta_{B}$
$A$ $A$ $B$
$B$ $C$ $A$
$C$ $B$ $C$
Table 6. (Top) A profile in which the plurality loser earns a small proportion of first-place votes and the election demonstrates a (runner-up) reinforcement paradox. (Bottom) A partition of the ballots which demonstrates the reinforcement paradox.
@@
To see that that bounds are tight, consider the profile in Table 7 where $\epsilon_{A},\epsilon_{B},\epsilon_{1}$ and $\epsilon_{2}>0$ are arbitrarily small parameters. If $\epsilon_{B}>\epsilon_{2}-\epsilon_{A}$ then this profile satisfies the conditions in (2) of Proposition 1
@@
Plurality loser case. From Proposition 1, $Z>\frac{1}{3}(X+Y)$ and thus the plurality loser must have at least $25\%$ of the first-place vote. On the other hand, consider the profile in Table 8 in which the plurality loser has just over 25% of the vote and yet, for some suitably small $\epsilon>0,$ satisfies the plurality loser inequalities from Proposition 1.

### v2

@@ Abstract.
We analyze the susceptibility of instant runoff voting (IRV) to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two sets of ballots such that a given losing candidate wins each of the sub-elections. Using these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We also analyze the frequency with which the paradox occurs in a large dataset of real-world ranked-choice elections to provide empirical probabilities. Our general finding is that IRV is highly susceptible to this paradox in three-candidate elections.
@@ 1. Introduction
The third strand is the expanding empirical social choice literature on IRV, enabled by the increasing availability of ranked-choice ballot data from jurisdictions such as Australia [43], the United States [34], and Scotland [24]. Some of this literature concerns the empirical probability that IRV produces a particular kind of paradox. Graham-Squire and Zayatz [15], Graham-Squire and McCune [17], and McCune and Graham-Squire [24] estimate the frequency with which real-world IRV elections demonstrate monotonicity and no-show paradoxes. Other empirical work includes case studies of particularly interesting elections which demonstrate a range of different paradoxes; see [20, 25], for example. Beyond work on paradoxes, empirical research on IRV has analyzed rates of ballot errors in IRV elections [7, 35], the extent of ballot exhaustion [6, 17], the rate at which IRV exhibits the so-called “spoiler effect” [27], the extent to which ballot truncation can affect the IRV winner [9, 19, 44], and the frequency with which IRV selects the Condorcet winner [17, 42, 43]. There is also a growing, related empirical literature concerning the method of the single transferable vote (STV), a multiwinner version of IRV which is designed to achieve proportional representation given a set of preference ballots. This literature also investigates paradoxical and Condorcet-related outcomes [24, 26], as well as STV’s propensity to achieve proportional representation [1, 3, 30].
@@
Previous work on the reinforcement paradox is fairly limited. The notion of a reinforcement paradox dates to the work of Young [45, 46], who defined the concept and proved essentially that the only reasonable voting methods not susceptible to the paradox are positional scoring rules. Felsenthal and Nurmi [13] show that various Condorcet inconsistent voting rules (including IRV) are susceptible to the paradox even in “tame” profiles in which a Condorcet winner exists and the voting rule chooses this candidate. Brandt et al. [2] provide some “minimal” examples of the reinforcement paradox under IRV where the number of candidates and voters is small. Brandt et al. [4] analyze the frequency of reinforcement paradoxes under various Condorcet consistent voting rules. We are aware of only two previous papers which attempt to measure the frequency with which IRV produces reinforcement paradoxes, both of which focus on the three-candidate case. Plassmann and Tideman [37] measure this frequency under a spatial model, and Courtin et al. [8] use the impartial culture and impartial anonymous culture models of voter behavior. Our work differs from these studies in significant aspects which we explicate more fully below. In brief, these two studies tend to generate two elections and then say that the reinforcement paradox has occurred if these two elections have the same winner but when the ballots from the two elections are combined the winner is different. By contrast, we consider only one election and say the paradox has occurred in this election if there exists a way to partition the ballots into two sub-elections where the winners of the sub-elections are the same but differ from the winner of the original election.
@@
The paper is structured as follows. Section 2 gives preliminaries, including formal definitions of IRV and the reinforcement paradox. In Section 3 we provide necessary and sufficient conditions for a three-candidate election to demonstrate a reinforcement paradox and discuss some of the consequences of these conditions. Proposition 1 gives these conditions assuming all voters provide a complete ranking of the candidates while Proposition 2 gives conditions when some voters cast partial ballots. Section 4 considers the minimum percentage a losing candidate can earn and yet the election demonstrates a reinforcement paradox. Section 5 estimates the probability that an election demonstrates a paradox under various theoretical models of voter behavior. Section 6 analyzes real-world IRV elections from the United State and Scotland. This section provides empirical probabilities using a large dataset of elections, and provides a probabilistic analysis using non-parametric bootstrapping. Section 7 gives a discussion of our results, and Section 8 concludes.
@@ 3. Necessary and Sufficient Conditions
The purpose of this article is to investigate the probability with which a three-candidate election demonstrates a reinforcement paradox. To examine such probabilities, we first must know when an election demonstrates a paradox. This section characterizes when an election demonstrates a reinforcement paradox, identifying sets of inequalities in terms of the parameters provided in Table 3 for the case of complete ballots or the parameters provided in Table 4 for the partial ballot case.
First, we observe that if one of the candidates earns an initial majority of first-place votes (we term such a candidate a “majority candidate”) then the election cannot demonstrate a reinforcement paradox. That is, we cannot observe a paradox in the complete ballot case if $X_{1}+X_{2}>Y_{1}+Y_{2}+Z_{1}+Z_{2}$ or $Y_{1}+Y_{2}>X_{1}+X_{2}+Z_{1}+Z_{2}$ or $Z_{1}+Z_{2}>X_{1}+X_{2}+Y_{1}+Y_{2}$; similarly for the partial ballot case. The reason is that if a majority candidate exists, that candidate will earn a majority of first-place votes in either $\mathcal{B}_{1}$ or $\mathcal{B}_{2}$ for any partition of the ballot set.
@@
The conditions in Proposition 2 reduce to those in Proposition 1 when $X_{3}=Y_{3}=Z_{3}=0$ for both runner-up and plurality loser reinforcement paradoxes. The third condition $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})$ in Case (2) of Proposition 2 reduces to $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})$ and does not have an analogue in Proposition 1 because it is implied by the other conditions in Case (2). To see this, note that
@@
Before sketching the proofs of the propositions, we make a few observations. Firstly, a runner-up and plurality loser reinforcement paradoxes can occur when the IRV and plurality winners are distinct or when they coincide, assuming either complete or partial ballots. Secondly, in Case (2) of the runner-up paradox and in the plurality loser paradox it is easily seen that the runner-up (plurality loser respectively) wins both sub-elections only when different candidates are eliminated in the two sub-elections.
Thirdly, in Case (1) of the runner-up paradox, the runner-up must be the plurality winner, and the proof of the propositions show that if the two conditions are met, the runner-up wins a majority of first-place votes in one sub-election and wins in two rounds with the IRV winner being the plurality loser in the other sub-election. Note in Case (1), if $X_{1}>X_{2}$ then $Y^{\prime}>-X_{1}+X_{2}+Z^{\prime}$ is automatically satisfied. Thus, if among the voters whose first choice is the IRV winner, more voters second-rank the runner-up than second-rank the plurality loser, the paradox occurs whenever the runner up is the plurality winner ($Y^{\prime}>X^{\prime}$ or $Y>X$ in the complete ballot case).
@@
Candidate $C$ cannot have a majority of first-place votes in both sub-elections (otherwise it would have a majority in the original election). We claim that $C$ cannot have a majority of first-place votes in either sub-election. To see this, suppose $C$ has a majority in $\mathcal{B}_{1}$. If $A$ is the plurality loser in $\mathcal{B}_{2}$, then
@@
Thus, $C$ must win in the second round after the elimination of a plurality loser in both sub-elections. The two sub-elections cannot have the same plurality loser (otherwise that candidate would be a plurality loser when the ballots were combined). So assume $B$ is the plurality loser in $\mathcal{B}_{1}$ and $A$ is the plurality loser in $\mathcal{B}_{2}.$ This yields the following set of inequalities.
@@
The inequalities created in each step of the iterative process in the proofs of the propositions can be used in reverse to identify election partitions leading to the paradoxes. For instance, from the proof of Case (1) of Proposition 1, we can derive the following bounds for the sub-elections.
@@
Care must be taken when applying these bounds: the inequalities hold only in the limit as no allowance is made for the discrete mature of ballot counts and the possibility of ties. However, if the $X_{1},X_{2},$ etc. values in the original election are sufficiently distinct, these bounds can be used to identify two sub-elections.
For instance, in the complete ballot case, if $X_{1}=100,X_{2}=105,Y=215,Z_{1}=110$ and $Z_{2}=90$, these conditions allow us to select (among other possibilities)
$x_{1}=99,x_{2}=1,y=106\mbox{ and }z=197,$
which leads to
@@
3.1. Consequences of the Algebraic Conditions
The inequalities identified in Propositions 1 and 2 allow us to identify further anomalies. For instance, Table 5 demonstrates that it is possible for a candidate to be the IRV winner, a Condorcet winner and a plurality winner and yet the profile still satisfies the requirements in Propositions 1 for both a runner-up and plurality loser paradox.
Num. Voters 1001 50 521 480 561 439
1st $A$ $A$ $B$ $B$ $C$ $C$
2nd $B$ $C$ $A$ $C$ $A$ $B$
3rd $C$ $B$ $C$ $A$ $B$ $A$
Table 5. (Top) A profile with the IRV, plurality, and Condorcet winners coinciding that demonstrates both runner-up and plurality loser paradoxes.
The next example shows that reinforcement paradoxes can be nested: an election may exhibit a runner-up or plurality loser paradox and each sub-election demonstrate an additional reinforcement paradoxes.
Example 3.
Consider the profile in Table 6 where $A$ is the IRV winner. The election demonstrates a runner-up paradox and $B$ wins in both sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. However each of these sub-elections themselves satisfies the conditions for a runner-up paradox. This means that sub-election $\mathcal{B}_{1}$, in which $C$ is the plurality loser and $A$ is the runner-up, can be further subdivided into two smaller sub-elections in which $A$ is the IRV winner. (One such partition is $\mathcal{B}_{11}$ which corresponds to $X_{1}=12$, $X_{2}=31$, $Y_{1}=0$, $Y_{2}=25$, $Z_{1}=0$ and $Z_{2}=17$ and $\mathcal{B}_{12}$ which corresponds to $X_{1}=28$, $X_{2}=31$, $Y_{1}=5$,$Y_{2}=26$, $Z_{1}=1$ and $Z_{2}=31$.) And sub-election $\mathcal{B}_{2}$, in which $A$ is the plurality loser and $C$ is the runner-up, can be further subdivided into two smaller sub-elections in which $C$ is the IRV winner. (One such partition is $\mathcal{B}_{21}$ which corresponds to $X_{1}=50$, $X_{2}=0$, $Y_{1}=4$, $Y_{2}=25$, $Z_{1}=30$ and $Z_{2}=0$ and $\mathcal{B}_{22}$ which corresponds to $X_{1}=10$, $X_{2}=38$, $Y_{1}=41$, $Y_{2}=54$, $Z_{1}=69$ and $Z_{2}=2$.)
Num. Voters 100 100 50 130 100 50
1st choice $A$ $A$ $B$ $B$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
Partition $\mathcal{B}_{1}$ 40 62 5 51 1 48
$\mathcal{B}_{2}$ 60 38 45 79 99 2
Table 6. A profile that demonstrates a runner-up paradox with sub-elections that demonstrate additional reinforcement paradoxes.
Additional instances of the paradox may occur in which an election can be divided into three or more sub-elections in which either the runner-up or plurality loser wins.
@@ 3.1. Consequences of the Algebraic Conditions
1st choice $A$ $B$ $C$ $C$
2nd choice $C$ $C$ $A$ $B$
3rd choice $B$ $A$ $B$ $A$
Table 7. A profile that demonstrates a plurality loser, but not a runner-up, reinforcement paradox.
Finally, we note that elections may be susceptible to a plurality loser paradox but not a runn-erup paradox, as demonstrated by the profile in Table 7 where neither Condition (1) nor (2) of Proposition 1 is satisfied, but the plurality loser inequalities are satisfied. Such preference profiles can arise when voter preferences are “single-peaked.” In this case, candidate $C$ is like a “centrist” candidate in a spatial model, as if $A$, $B$, and $C$ have been arranged on a left-to-right line, with $C$ in the middle.
Of course, real-world elections typically do not return ballot data resembling the profile in Table 7. In fact as we will see in Section 6 in real elections, the IRV winner tends to be a much stronger candidate than the plurality loser. Thus we expect that if an election demonstrates a plurality loser reinforcement paradox in this data then the election will also demonstrate a runner-up paradox.
@@ 4. Threshold Results
Consider the profile in Table 8, where $p_{A}$ is the proportion of voters who cast the ballot $A\succ B\succ C$ and also the proportion of voters who cast the ballot $A\succ C\succ B$, and $\epsilon$ and $\delta$ are positive and very small. Furthermore, assume $0.52p_{A}+\epsilon<$ 0.5 and $14p_{A}+\epsilon+\delta=$ 1 (which implies $\delta>\epsilon)$. Note that $A$ is the IRV winner and $C$ is the plurality loser.
This profile satisfies condition (2) of Proposition 1 and thus the election demonstrates a runner-up reinforcement paradox. For completeness, we include a partition of the electorate into two sub-electorates, as shown in the table. Let $\delta_{B}$ and $\delta_{A}$ be positive small parameters such that $20<\delta_{B}<\delta_{A}<\delta/$ 2. Then $B$ wins when using the ballots of both sub-electorates separately. Since $\delta$ is arbitrary, the proposition is proven.∎
Prop. Voters $p_{A}$ $p_{A}$ $2p_{A}+\epsilon$ $\delta$
1st choice $A$ $A$ $B$ $C$
2nd choice $B$ $C$ $A$ $A$
3rd choice $C$ $B$ $C$ $B$
Partition $\mathcal{B}_{1}$ $\delta-\delta_{A}$ 0 $\delta-\delta_{B}$ $\delta$
$\mathcal{B}_{2}$ $p_{A}-\delta+\delta_{A}$ $p_{A}$ $2p_{A}+\epsilon-\delta+\delta_{B}$ 0
Table 8. A profile in which the plurality loser earns a small proportion of first-place votes and the election demonstrates a (runner-up) reinforcement paradox.
@@
To see that that bounds are tight, consider the profile in Table 9 where $\epsilon_{A},\epsilon_{B},\epsilon_{1}$ and $\epsilon_{2}>0$ are arbitrarily small parameters. If $\epsilon_{B}>\epsilon_{2}-\epsilon_{A}$ then this profile satisfies the conditions in (2) of Proposition 1.
@@
Plurality loser case. From Proposition 1, $Z>\frac{1}{3}(X+Y)$ and thus the plurality loser must have at least $25\%$ of the first-place vote. On the other hand, consider the profile in Table 10 in which the plurality loser has just over 25% of the vote and yet, for some suitably small $\epsilon>0,$ satisfies the plurality loser inequalities from Proposition 1. ∎
@@
We conclude this section by noting that there are elections that demonstrate a reinforcement paradox in which one of the sub-elections is arbitrarily small.
Example 4.
Let $m\in\mathbb{N}$ be a large even number. Consider an election with ballot set $\mathcal{B}$ where $m$ ballots have the form $A\succ B$, $m+1$ have the form $B\succ A$, $m/2$ have the form $C\succ A$, and $m/2-2$ have the form $C\succ B$. Then $A$ wins this election, since initially they have one fewer vote than $B$ but the elimination of $C$ allows $A$ to win by one vote.
Let $\mathcal{B}_{1}$ consist of two ballots of the form $A\succ B$, three of the form $B\succ A$, and four of the form $C\succ A$. Let $\mathcal{B}_{2}$ consist of $m-2$ ballots of the form $A\succ B$, $m-2$ of the form $B\succ A$, $m/2-4$ of the form $C\succ A$, and $m/2-2$ of the form $C\succ B$. Then $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ form a partition of $\mathcal{B}$, and candidate $B$ wins both of these sub-elections. Since $m$ can be arbitrarily large, this implies there exists an election demonstrating a reinforcement paradox where one of the sub-elections can be arbitrarily small.
Example 4 is interesting as it suggests that $B$ was “close” to being a winner in the original election. In fact, $B$ might feel it particularly unfair that the addition of a handful of ballots (from $\mathcal{B}_{1}$, where $B$ was a winner) caused $B$ to go from winner (in $\mathcal{B}_{2}$) to loser (in $\mathcal{B}_{1}\cup\mathcal{B}_{2}$).